UT 06 Sep 2021 - Night #5 of Restart




Night Staff:

OSA: S. Allanson

ISp: R. Reyes

ISA: B. Rothberg, R.T. Gatto (Tucson secret Outpost)

Technical Wizard: J. Hill (Tucson)

Mtn Manager: P. Hartley

Sunset: 18:43

Twlight Ends: 20:03

Twilght Starts: 04:34

Sunrise: 05:53

Starting Config: LUCI-Bino (for AGW 1/2)

Night Log:

UT 23hrs-ish: Power bump on mountain and VPN and some intenret services on campus went down - power and internet up on mountain - using mountain vpn for now.

UT 23:41 - starting up LUCI-1 GUIs, will init all and take test exposures

luci1.20210906.0001-0002 - 2.51sec test dark

-look OK

UT 01:22 - VPN to campus seems to be back, although did lose the connection to ROBS (both an ssh terminal AND the X2GO)

UT 01:40 - staying closed for now

UT 03:29 - sucker holes come and go, but there's a constant ring of dark clouds visible, humidity seems to be edging upwards.

UT 04:19 - internet slow - mountain to campus is maxed out at 4Mbps, but upstream is 200+ Mbps. No one on-call at UITS. Acessing stuff FROM the mountain is super slow (kb/s speeds on average).

image_2021_09_06T04_18_08_464Z.png

UT 04:54 -clouds thinning out, humidity down to 57%

-may do StoneP and BS9188 to get 16 data points in total for field aberration data - not sure next time we'll have a shot at this...

UT 05:06 - opening vent door....

UT 05:10 - opening enclosure

Field Aberration Data Collection Binocular Mode

Stone P1

UT 05:17 - moving to pointing star near Stone P1

UT 05:22 - collimating at Stone P1

UT 05:28 - NOW starting field_collect, too slow to ssh remote display X-windows from OBS2 to ROBS, 2nd X2GO now on OBS2 to run this

-set limit=350 nm RMS, but starting to suspect that keyword doesn't actually do anything...

starting airmass ~1.3

seeing generally below 1.2" mostly 0.8-1.1" so far

UT 05:53 - seeing blew up a bit to 1.6" in guider, might be a quick bubble

UT 05:56 -fuzzy PSFagain on guider

UT 06:02 - seeing getting progressively worse 1.9-2" on DX!

UT 06:08 - now back down to 1.2"

UT 06:11 - bow back up to...wait...was 1.5" now 1.1"

UT 06:15 - ended on 1.8"! in guider,

Not confident in how good this data is....

LFBG Field Results Stone_P1 (8 stars)
 => Field Data Results
      Motion of the secondary needed to remove bi-nodal astigmatism

                  Stone_P1
              From residual Focus Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  25.0", Delta IE = -25.0"
         Tilt =  -4.3", Delta CA =   4.3"


 => Field Data Results
             Motion of the secondary needed to remove bi-nodal astigmatism

                  Stone_P1
              From Total Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   0.2", Delta IE =  -0.2"
         Tilt =  -1.1", Delta CA =   1.1"


 => Field Data Results
             Measured bi-nodal astigmatism

                  Stone_P1
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  32.7", Delta IE = -32.7"
         Tilt = -45.3", Delta CA =  45.3"


                  Stone_P1
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  34.2", Delta IE = -34.2"
         Tilt = -32.3", Delta CA =  32.3"


                  Stone_P1
              From both Astigmatisms Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  36.7", Delta IE = -36.7"
         Tilt = -21.7", Delta CA =  21.7"

RFBG Field Results Stone_P1 (8 stars)
 => Field Data Results
      Motion of the secondary needed to remove bi-nodal astigmatism

                  Stone_P1
              From residual Focus Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   5.1", Delta IE =  -5.1"
         Tilt =   9.8", Delta CA =  -9.8"


 => Field Data Results
             Motion of the secondary needed to remove bi-nodal astigmatism

                  Stone_P1
              From Total Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   0.1", Delta IE =  -0.1"
         Tilt =   0.6", Delta CA =  -0.6"


 => Field Data Results
             Measured bi-nodal astigmatism

                  Stone_P1
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -23.1", Delta IE =  23.1"
         Tilt =  22.3", Delta CA = -22.3"


                  Stone_P1
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   3.5", Delta IE =  -3.5"
         Tilt =  19.3", Delta CA = -19.3"


                  Stone_P1
              From both Astigmatisms Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -17.3", Delta IE =  17.3"
         Tilt =  18.3", Delta CA = -18.3"

BS9188

UT 06:17 - pointing star near BS9188

UT 06:22 - moving BS9188 with active preset to check IQ there (airmass ~1.05 instead of 1.4-1.2)

UT 06:25 - Definitely looks better here at aimrass 1.05 - sending GUIDE preset then will execute more data collection

UT 06:26 - starting data collection

UT 07:10 - DX done

UT 07:11 - SX Done

LFBG Field Results BS9188 (8 stars)
 => Field Data Results
      Motion of the secondary needed to remove bi-nodal astigmatism

                  BS9188_field
              From residual Focus Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  -6.0", Delta IE =   6.0"
         Tilt = -15.5", Delta CA =  15.5"


 => Field Data Results
             Motion of the secondary needed to remove bi-nodal astigmatism

                  BS9188_field
              From Total Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   2.2", Delta IE =  -2.2"
         Tilt =  -3.9", Delta CA =   3.9"


 => Field Data Results
             Measured bi-nodal astigmatism

                  BS9188_field
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  -7.3", Delta IE =   7.3"
         Tilt = -47.9", Delta CA =  47.9"


                  BS9188_field
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   9.0", Delta IE =  -9.0"
         Tilt = -16.9", Delta CA =  16.9"


                  BS9188_field
              From both Astigmatisms Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   3.1", Delta IE =  -3.1"
         Tilt = -27.5", Delta CA =  27.5"

RFBG Field Results BS9188 (8 stars)
 => Field Data Results
      Motion of the secondary needed to remove bi-nodal astigmatism

                  BS9188_field
              From residual Focus Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -16.5", Delta IE =  16.5"
         Tilt =   1.0", Delta CA =  -1.0"


 => Field Data Results
             Motion of the secondary needed to remove bi-nodal astigmatism

                  BS9188_field
              From Total Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   2.6", Delta IE =  -2.6"
         Tilt =   3.0", Delta CA =  -3.0"


 => Field Data Results
             Measured bi-nodal astigmatism

                  BS9188_field
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -42.0", Delta IE =  42.0"
         Tilt = -29.2", Delta CA =  29.2"


                  BS9188_field
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  -6.7", Delta IE =   6.7"
         Tilt =   5.4", Delta CA =  -5.4"


                  BS9188_field
              From both Astigmatisms Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip = -27.7", Delta IE =  27.7"
         Tilt = -16.6", Delta CA =  16.6"

LFBG Field Combined Results (16 stars)
 => Field Data Results
      Motion of the secondary needed to remove bi-nodal astigmatism

                  TwoTarg
              From residual Focus Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  11.1", Delta IE = -11.1"
         Tilt =  -9.6", Delta CA =   9.6"


 => Field Data Results
             Motion of the secondary needed to remove bi-nodal astigmatism

                  TwoTarg
              From Total Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =   0.8", Delta IE =  -0.8"
         Tilt =  -2.3", Delta CA =   2.3"


 => Field Data Results
             Measured bi-nodal astigmatism

                  TwoTarg
              From 45 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  14.7", Delta IE = -14.7"
         Tilt = -46.6", Delta CA =  46.6"


                  TwoTarg
              From 0 Astigmatism Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  22.1", Delta IE = -22.1"
         Tilt = -25.5", Delta CA =  25.5"


                  TwoTarg
              From both Astigmatisms Measurement
                  On-Sky Arcseconds
         -------------------------------------------
          Tip =  21.3", Delta IE = -21.3"
         Tilt = -24.7", Delta CA =  24.7"

LFBG Field Corrections Applied (afternoon of 20210911)
Dear Telescope,
     I have made an update of the SX M2 positions to make a field
aberration correction.  To all Gregorian focal stations (except ARGOS)
I have adjusted M2 for Tip=+15 Tilt=-25.  This was an estimated correction
based on the BS9188 and Stone_P1 field data collected on 20210906 UT.

The corrections are presently kept in the Global Offsets buffer, so you
may notice them there on the PSFGUI.

X=+1.277 Y=-0.766 Z=0 RX=+147.3 RY=+245.5

We are now ready to take a confirmation set of field data (about 8 stars)
on either LFBG or LDG.  No changes were made to RFBG where the corrections
were small.
Further refinements were made after the confirmatory data.


LUCI-1 Checkout

UT 07:12 - authorize LUCI-1 only

-OBS2 was so much quicker than ROBS is right now...
Field Center Data Collection

Target: M52

UT 07:17 - putting N30 Fieldstop into FPU

UT 07:19 - sending Active Preset to M52 PA=0
  • L1 0003- active on
  • L1 0004 - background image
  • L1 0005 - Track PA 0
  • L1 0006 - Track 180
  • L1 0007 - Track 90
  • L1 0008 - Track 270

Rotator ZP Checks with Astrometric MOS Masks
Stone O

UT 07:33 - Rotator ZP Stone O Preset

#36 thought wrongs star selected, choosing other star makes it worse. Outliers are 34, 35, 36

Remove #36 impcoves, Remove #35 good fit

Err RA 0.070, Dec 0.079

xshift = -0.192", yshift=-2.105", theta=0.05

L1 0012 - thru-slit verification

Stone P

UT 07:50 sending preset to Stone P

Once again RTD not respnse, reset all, loaded images back

initial fit is mediocre in Dec, #36 worst outlier -removing it

Err RA 0.073, Dec 0.091

Xshift = -0.142', Yshift = -2.175" theta = 0.105

L1 0016 - thru slit verification
Stone A

UT 08:05 - sending preset to StoneA

pixel map won't load - reseting RTD

initial fit - Good for all points

Err= RA 0.067, Dec 0.052

Xshift = -0.646", Yshift = -2.321", theta=0.0464

Mean Theta = 0.067 deg, Std Dev = 0.032 deg

-not updating the Rotator zero-point - seems ok

AGW Transform LUCI-1

Target Stone A3 - using list Stone_A3_LUCI.list

UT 08:23 - pointing check

-putting N30 fieldstop into place w/N375 camera, H-filter DIT=4sec, flexure off

UT 08:32 - send Acquire preset to Stone A3

-forgot to take blind out, did it in time befor first exposure

-its 2.51 sec, it ignored my exposure time for DIT=4 (must be hard-coded) - ring of holes has 1800 counts - so ok
  • L1 0020 - Dit =4 blind filter - might not be useful
  • L1 0021 - sky image
  • Following are PA=180
  • L1 0023 - GS 10
  • L1 0024 - GS 5
  • L1 0025 - GS 1
  • GS 13 out of range - contintuing at next one
  • L1 0026 - GS 8
  • L1 0027 - GS 5 (repeated star at same PA as above)
  • L1 0028 - GS 4
  • L1 0029 - GS 3
  • L1 0030 - GS 7
  • GS 18 out of reach - (x,y): 271.808, -3.70891 mm.
  • GS 19 out of reach - (x,y): -13.3087, 275.056 mm.
  • Following are PA=170
  • GS 10 out of reach - (x,y): -183.371, 9.33718 mm
  • L1 0031 - GS 5
  • L1 0032 - GS 1
  • L1 0033 - GS 13
  • L1 0034 - GS 8
  • L1 0035 - GS 5 (repeated star at same PA as above)
  • L1 0036 - GS 4
  • L1 0037 - GS 3
  • L1 0038 - GS 7
  • GS 18 out of reach - (x,y): 268.344, 43.5535 mm
  • GS 19 out of reach - (x,y): -60.8475, 268.571 mm
  • L1 0039 - GS 9
  • L1 0040 - GS 0
  • Removed the out of range one in file and added GS 9 170 at end
  • 16 stars obtained
  • data/20210906/lfbg/transform_luci1.log
  • Seeing <1" for all observations
-somehow clicked a mouse click that GEIRS only lets me change stretch and nothing else

Imaging Tests
Dither out of field - Sidereal

Target NGC 40 - cycle through some filters

UT 09:20 - putting N30 field stop mask away

UT 09:21 - restarting the GEIRS window

UT 09:23 - sending preset to NGC 40 - bit of a slew from Stone A3
  • L1 0041-0082
  • Full K,H,J, 4 exp of H2, Full HeI
Dither out of Field -Non-Sidereal

NSIGUI on ROBS doesn't seem to work , same error on OBS2 also on OBS1 - IT 8462

[lbto@robs ~]$ can't read "ftpName": no such variable
while executing
"send "get $ftpName $localFile\r""
invoked from within
"expect "ftp> " {send "get $ftpName $localFile\r"}"
(file "/usr/bin/nsQueryEngine.tcl" line 361)

It appears to retrieve the item with the ephemeris in a .log file, but the .txt (whch is read in) has 0 counts.

UT 10:37 - sending preset - I had already generated the ephemeris for UT Sep 06 - so using that file

-testing big offsets and return to field with NS motion
  • What on earth... the probe is on-axis ... ah I think its because I accidentally clicked Non-SIdereal Guide override. Barry mistake (click happy)
  • L1 0082-0108


Photometric Standards

UT 11:30 - sending preset to P161D - photometric standard - zJHKKs - all-sky looks fairly clear
  • Airmass 1.39
  • L1 0109-0128
UT 12:05 - sending preset to P247U
  • Airmass 1.005
  • only 3 positions for H-band, only 3 positions for J-band
  • L1 0129-0146

UT 12:32 - closing Dome

LUCI-1 Imaging Cals

  • z-band L1 0147-0156
  • J-band L1 0157-0166
  • HeI L1 0167-0176
  • H-band L1 0177-0186
  • K-band L1 0187-0196
  • Ks-band L1 0197-0206
UT 13:24 - putting LUCI-1 in safe config

Proposed Plan of Action:

  • LUCIs (AGWs):
    • DONE Field Aberration Data Collection
      • Fields:
      • BS9177 18:39:33.7 +49:05:38.0
      • DONE StoneP1 22:37:45.0 -01:13:40.45
      • DONE BS9188 23:30:33.4 +38:18:57.0
      • BS9107 03:32:03.0 +37:20:40.0
  • LUCIs
    • DONE Field Center Rotation Check (0.5hr) - one field of the three
      • M92 - start
      • M34 - mid
      • M52 - end
    • DONE measure rotator zeropoint (requires MOS mask) - (0.5 hr)
      • ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
    • DONE AGW Transform (if seeing < 1.2")
      • StoneO 21 37 31 +00 31 55 Stone_O_m12.list
      • StoneP 22 35 49 +01 31 33 Stone_P1.list
      • DONE StoneA 00 54 02 -00 32 34 Stone_A3.list
    • Imaging check
      • dither in field (0.5hr) -
      • DONE dither out of field (0.5hr) - NGC 040
    • non-sidereal Imaging Check
      • dither in field
      • DONE dither out of field - Comet 69P CG
        • Comet 69/P CG
  • LBCs (if clear any seeing)
    • Standard photometric fields (new ones ready) - requires clear skies (any seeing) - required for ZP calibrations
    • Spectro-photometric standards - requires clear skies (any seeing) - required for ZP calibrations
    • Twilight Flats all filters - High (4 x 0.5 hrs - likely min two nights to complete)
    • Poor Conditions
      • LBC Pointing Model - (3hr) - any conditions
        • Standard Method
          • <a href='https://wiki.lbto.org/bin/edit/Commissioning/PtStarrs?topicparent=Commissioning.Com2021Restart' rel='nofollow' title='Create this topic'> PtStarrs </a> - please choose targets > 80 deg elevation (more than 1 or 2)
          • Run TMS passively - run 1 reference at the beginning
          • Need stable temperatures
          • Turn on PCS logging
          • Need to collimate at each star
          • Chip 2 only exposure

Current Status

  • LBCs
    • LBC-Blue - available
    • LBC-Red - available
  • MODS
    • MODS-1 - offline - (computers OK, but not ready to pump until Monday)
      • AGW-5 - available
    • MODS-2 - available
      • AGW-6 - available
  • LUCI-1 - available
    • AGW-1 available
  • LUCI-2 offline
    • AGW-2 available
  • PEPSI - offline Software environmental system issues - earliest Monday
    • AGW-7/8 available


Closed Dome Activities

  • ARGOS Cal-Unit test for M3 alignment and pupil wobble checkout (alternate to on-sky method)
  • LUCI-1 Only Calibrations
    • Flexure data for G150
      • /home/dthomspons/Luci/FlexureCompensation
      • README - Calib15 - 2hrs
      • Ks filter, Argon lamp, optical sieve mask
      • take dark frame
      • Manual Control MCSGUI for elevations
    • Wavelength data for G150
      • G150 Ks, Argon, take image through slit, then put in grating, line should show up where the slit is located (within a few pixels of the image of the slit)
      • Check a few lines (2 or 3)

-- BarryRothberg - 05 Sep 2021
Topic revision: r8 - 17 Sep 2021, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback